OpenStudy (anonymous):

complete the square to find the center and radius of the circle whose equation is x^2 + y^2 + 8x - 4y = 36

OpenStudy (anonymous):

@dan815

OpenStudy (campbell_st):

you need to complete the square in both x and y \[(x^2 + 8x+ ?) + (y^2 -4y+ ??) = 36\] the values you add to complete the square need to be added to both sides of the equation. then factor the standard form is \[(x -h)^2 + (y - k)^2 = r^2\] where (h, k) is the centre and r is the radius of the circle hope it helps

OpenStudy (anonymous):

how do i find h k and r

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so do you know how to complete the square..?

OpenStudy (anonymous):

yeah four and -2 work

OpenStudy (campbell_st):

ok... so in x it becomes \[(x^2 + 8x + 16) + (y^2 - 4y + 4) = 36 + 16 + 4\] the values you add to the left, to complete the square need to be added to the right. so next step is to factor the perfect squares and simplify the right side what would you get..?

OpenStudy (anonymous):

(x+4)(x+4) and (y-2)(y-2)

OpenStudy (anonymous):

=56

OpenStudy (campbell_st):

that's ok... the more correct way to write it is \[(x + 4)^2 + (y - 2)^2 = 56\] so if you compare it to the standard form \[(x -h)^2 + (y - k)^2 = r^2\] you should be able to identify the centre and radius

OpenStudy (anonymous):

thank you

OpenStudy (campbell_st):

(h, k) is the centre and r is the radius

OpenStudy (anonymous):

so (4,-2) and 56